OpenStudy (anonymous):

Jennifer found some boxes of clothes in her grandmother’s attic. She decided to donate 3/8 of the boxes to a local charity. What fraction is 3/8 equivalent to ?

OpenStudy (anonymous):

PLEASE HELP!!

OpenStudy (anonymous):

be more specific?? 3/8 is equivalent to a lot of things. 3/8 = 6/16 = 9/24 = 30/80 = 375/1000 = etc.

OpenStudy (anonymous):

Here are the opptions

OpenStudy (anonymous):

multiply the top and bottom by any number that you want (except zero), and you have an equivalent fraction

OpenStudy (anonymous):

9/32 12/24 6/16 18/64